OpenStudy (anonymous):

help!!!!! -x+y=12 x+2y=3

OpenStudy (agent47):

Add the equations: 0+3y=15 y=5

OpenStudy (agent47):

-x+5=12 -x=12-5 x=-7

OpenStudy (anonymous):

thanks

OpenStudy (unklerhaukus):

to check your answer \[-(-7)+(5)=12\] \[(-7)+2\times(5)=3\]

OpenStudy (anonymous):

system of inequality y less than equal to +3 y greater = to -x-2

OpenStudy (unklerhaukus):

\[y≤+3\]\[y≥-x-2\]

OpenStudy (anonymous):

yes

OpenStudy (unklerhaukus):